# NeRF-LiDAR: Generating Realistic LiDAR Point Clouds with Neural Radiance Fields

### 09 Jan 2024 v2

@@ Abstract
Labelling LiDAR point clouds for training autonomous driving is extremely expensive and difficult. LiDAR simulation aims at generating realistic LiDAR data with labels for training and verifying self-driving algorithms more efficiently. Recently, Neural Radiance Fields (NeRF) have been proposed for novel view synthesis using implicit reconstruction of 3D scenes. Inspired by this, we present NeRF-LIDAR, a novel LiDAR simulation method that leverages real-world information to generate realistic LIDAR point clouds. Different from existing LiDAR simulators, we use real images and point cloud data collected by self-driving cars to learn the 3D scene representation, point cloud generation and label rendering. We verify the effectiveness of our NeRF-LiDAR by training different 3D segmentation models on the generated LiDAR point clouds. It reveals that the trained models are able to achieve similar accuracy when compared with the same model trained on the real LiDAR data. Besides, the generated data is capable of boosting the accuracy through pre-training which helps reduce the requirements of the real labeled data.
@@ Introduction
State-of-the-art LiDAR simulation methods (Fang et al. 2020; Manivasagam et al. 2020) first reconstruct the real-world driving scenes into 3D meshes and then run the physics-based simulation. In order to achieve dense accurate reconstruction results, these methods need to scan the street many times using expensive LiDAR devices (Fang et al. 2020). More importantly, it’s still expensive to generate point-wise semantic labels for simulated LiDAR data as it requires human annotations on the reconstructed 3D scenes.
Instead of simulating the whole LiDAR scenes, others, e.g.,  (Fang et al. 2021) use the real-world 3D scenes and propose a rendering-based LiDAR augmentation framework to enrich the training data and boost the performance of LiDAR-based 3D object detection.
Our method also leverages real-world information for learning LiDAR simulation. Our NeRF-LiDAR creates an implicit neural-radiance-field representation of the real world for both point clouds and label rendering.
@@
Many recent NeRFs have been proposed to address the challenges of large-scale outdoor scenes  (Barron et al. 2021, 2022; Tancik et al. 2022; Zhang et al. 2020).
There are also some methods  (Rematas et al. 2022; Deng et al. 2022) leveraging depth supervision to help create more accurate 3D geometry of scenes. Panoptic or semantic label synthesis for novel views is also explored in (Zhi et al. 2021; Kundu et al. 2022; Fu et al. 2022). They utilize the density of the volume to render image labels along with the novel view synthesis.
Inspired by these works, our method reconstructs the accurate 3D geometry using the NeRF methodology in the driving scene and generates 3D point clouds along with accurate semantic labels for the LiDAR simulation.
@@
$\hat{\mathbf{C}}(\textrm{r})=\sum_{i=1}^{N}T_{i}(1-e^{-\sigma_{i}\delta_{i}})\mathbf{c}_{i},\quad T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right),$ (1)
@@
Different from them, we present a new method that can take multi-view images and sparse LiDAR signals to reconstruct the street scenes and represent the 3D scenes as an implicit NeRF model. Recently, more and more driving-scene data are collected by self-driving cars. We propose to use these data to learn the NeRF reconstruction of the driving scenes.
NeRF reconstruction of the unbounded large-scale driving scenes is challenging. This is because most of NeRFs (Mildenhall et al. 2020) are designed for small scene reconstruction with object-centric camera views. However, the driving data are often collected in the unbounded outdoor scenes without object-centric camera views settings (e.g., nuScenes (Caesar et al. 2020) dataset). Moreover, since the ego car moves fast during the data collection, the overlaps between adjacent camera views are too small to be effective for building multi-view geometry. We reconstruct the NeRF representation based on the multi-view images collected from the ego vehicle and leverage the LiDAR points to provide extra depth supervision to create more accurate 3D geometries. Besides, the real LiDAR point clouds are used as supervision to learn more realistic simulated LiDAR data.
To reconstruct the NeRF for driving scenes, we use the unbounded NeRF (Barron et al. 2022) with a modified supervision of:
@@
To achieve the point-wise semantic labels of the simulated LiDAR point clouds, we use the 2D semantic labels of the images to learn the 3D label generation.
@@
where $CE$ is the cross-entrophy loss, $\mathbf{r}$ represents pixel-wise camera rays corresponding to each image pixel, $\hat{\mathbf{S}(\mathbf{r}}$ is the rendered labels by the NeRF model (Eq.(8)) and $\mathbf{S}(\mathbf{r}$ is the label predicted by the image segmentation model.
@@
To project the irregular LiDAR points, we adopt the spherical projection  (Geiger, Lenz, and Urtasun 2012) to project our points into the equirectangular image grids (as illustrated in Figure 3).
Similarly, the real LiDAR frame is also transferred into a 2D $32\times 1024$ equirectangular image. In this way, we can easily create the correspondence in 2D grids.
@@
$\vspace{-3mm}\mathcal{L}_{feat}=\sum_{k=1}^{n}w_{k}\|F(\hat{I})_{k}-F(I_{gt})_{k}\|_{1},$ (14)
where $F$ is a pre-trained and fixed feature extractor (e.g., VGG (Simonyan and Zisserman 2015), Point-cloud segmentation Net (Milioto et al. 2019)). We use $n$-level pyramidal features to compute the feature distance. $n=4$ is the number of feature levels, $w_{k}=2^{k-n}$ is the weights for $k$th-level features. The loss $L_{feat}$ measures the feature-level similarity between the real and the simulated point clouds. To enable the back-propagation from the feature loss to the previous raydrop module, we apply the gumble-softmax proposed in (Jang, Gu, and Poole 2016) on the ray drop processing.
@@
no ray drop random drop learning drop vgg loss rangenet loss
Simulation 65.7
@@
Comparisons with CARLA and real LiDAR
In Table 1, we evaluate the quality of the generated data by using it to train 3D segmentation model (Zhou et al. 2021), the mean IoU is used as the evaluation metric.
@@
Comparisons with other reconstruction-based simulators
Our NeRF-LiDAR possesses apparent advantages over other reconstruction-based LiDAR simulators (Manivasagam et al. 2020; Fang et al. 2020). First, RGB images are used to assist the reconstruction in our NeRF-LiDAR, providing useful multi-view geometry information for more accurate reconstruction and label generation. Secondly, no manual annotations on the point clouds are required. We use NeRF representation to learn label generation. Multi-view images provide useful label information and geometry consistency to reduce the outlier labels. Moreover, no mesh reconstruction of the real world is required. NeRF-LiDAR can preserve more details of the real world by leveraging the RGB texture contents. Finally, NeRF-LiDAR does not require dense point clouds for the reconstruction of the real world. LiDARSim (Manivasagam et al. 2020) uses a 64-channel LiDAR and scans the street many times to achieve the dense point clouds for LiDAR simulation. Augmented LiDAR simulator (Fang et al. 2020) utilizes a special and expensive LiDAR device to achieve the dense point clouds. As a comparison, NeRF-LiDAR relies more on the 2D images to achieve 3D geometry information.
@@
Combining real and simulation data
We combine real data with NeRF-LiDAR data generated from ground-truth scenes to see if simulated data can further boost performance when used for training. As shown in Table 1, simulated data along with 10% real data (real 1k) are able to achieve better performance (Sim10k + real 1k: 89.1) than 100% real data (real 10k: 88.4).
@@ Appendix A More Experimental Results And Details
Table 5: Evaluation on the Test Set 1 (the nuScenes validation set), where the simulation data and the test set is from the same driving scenes. Rangenet++(Milioto et al. 2019) is trained on the different training sets for evaluations. $*{}^{*}$ mean IoU of four classes.

### 20 Jan 2024 v3

@@ Abstract
Labelling LiDAR point clouds for training autonomous driving is extremely expensive and difficult. LiDAR simulation aims at generating realistic LiDAR data with labels for training and verifying self-driving algorithms more efficiently. Recently, Neural Radiance Fields (NeRF) have been proposed for novel view synthesis using implicit reconstruction of 3D scenes. Inspired by this, we present NeRF-LIDAR, a novel LiDAR simulation method that leverages real-world information to generate realistic LIDAR point clouds. Different from existing LiDAR simulators, we use real images and point cloud data collected by self-driving cars to learn the 3D scene representation, point cloud generation and label rendering. We verify the effectiveness of our NeRF-LiDAR by training different 3D segmentation models on the generated LiDAR point clouds. It reveals that the trained models are able to achieve similar accuracy when compared with the same model trained on the real LiDAR data. Besides, the generated data is capable of boosting the accuracy through pre-training which helps reduce the requirements of the real labeled data. Code is available at https://github.com/fudan-zvg/NeRF-LiDAR
@@ Introduction
State-of-the-art LiDAR simulation methods (Fang et al. 2020; Manivasagam et al. 2020) first reconstruct the real-world driving scenes into 3D meshes and then run the physics-based simulation. In order to achieve dense accurate reconstruction results, these methods need to scan the street many times using expensive LiDAR devices (Fang et al. 2020). More importantly, it’s still expensive to generate point-wise semantic labels for simulated LiDAR data as it requires human annotations on the reconstructed 3D scenes. Instead of simulating the whole LiDAR scenes, others, e.g.,  (Fang et al. 2021) use the real-world 3D scenes and propose a rendering-based LiDAR augmentation framework to enrich the training data and boost the performance of LiDAR-based 3D object detection. Our method also leverages real-world information for learning LiDAR simulation. Our NeRF-LiDAR creates an implicit neural-radiance-field representation of the real world for both point clouds and label rendering.
@@
Many recent NeRFs have been proposed to address the challenges of large-scale outdoor scenes  (Barron et al. 2021, 2022; Tancik et al. 2022; Zhang et al. 2020). There are also some methods  (Rematas et al. 2022; Deng et al. 2022) leveraging depth supervision to help create more accurate 3D geometry of scenes. Panoptic or semantic label synthesis for novel views is also explored in (Zhi et al. 2021; Kundu et al. 2022; Fu et al. 2022). They utilize the density of the volume to render image labels along with the novel view synthesis. Inspired by these works, our method reconstructs the accurate 3D geometry using the NeRF methodology in the driving scene and generates 3D point clouds along with accurate semantic labels for the LiDAR simulation.
@@
$\hat{\mathbf{C}}(\textrm{r})=\sum_{i=1}^{N}T_{i}(1-e^{-\sigma_{i}\delta_{i}})\mathbf{c}_{i},\ T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right),$ (1)
@@
In this paper, we present a new method takes multi-view images and sparse LiDAR signals to reconstruct the street scenes and represent the 3D scenes as an implicit NeRF model. We propose to use the driving-scene data to learn the NeRF reconstruction.
NeRF reconstruction of the unbounded large-scale driving scenes is challenging. This is because most of NeRFs (Mildenhall et al. 2020) are designed for small scene reconstruction with object-centric camera views. However, the driving data are often collected in the unbounded outdoor scenes without object-centric camera views settings (e.g., nuScenes (Caesar et al. 2020)). Moreover, since the ego car moves fast during the data collection, the overlaps between adjacent camera views are too small to be effective for building multi-view geometry. We reconstruct the NeRF representation based on the multi-view images and leverage the LiDAR points to provide extra depth supervision to create more accurate 3D geometries. Besides, the real LiDAR point clouds are used as supervision to learn more realistic simulated LiDAR data.
To reconstruct the driving scenes, we use the unbounded NeRF (Barron et al. 2022) with a modified supervision of:
@@
To achieve the point-wise semantic labels of the simulated LiDAR points, we use the 2D semantic labels of the images to learn the 3D label generation.
@@
where $CE$ is the cross-entrophy loss, $\mathbf{r}$ represents pixel-wise camera rays corresponding to each image pixel, $\hat{\mathbf{S}}(\mathbf{r})$ is the rendered labels by the NeRF model (Eq.(8)) and $\mathbf{S}(\mathbf{r})$ is the label predicted by the image segmentation model.
@@
To project the irregular LiDAR points, we adopt the spherical projection  (Geiger, Lenz, and Urtasun 2012) to project our points into the equirectangular image grids (as illustrated in Figure 3). Similarly, the real LiDAR frame is also transferred into a 2D $32\times 1024$ equirectangular image. In this way, we can easily create the correspondence in 2D grids.
@@
$\mathcal{L}_{feat}=\sum_{k=1}^{n}w_{k}\|F(\hat{I})_{k}-F(I_{gt})_{k}\|_{1},$ (14)
where $F$ is a pre-trained and fixed feature extractor (e.g., VGG (Simonyan and Zisserman 2015), Point-cloud segmentation Net (Milioto et al. 2019)). We use $n$-level pyramidal features to compute the feature distance. $n=4$ is the number of feature levels, $w_{k}=2^{k-n}$ is the weights for $k$th-level features. The loss $L_{feat}$ measures the feature-level similarity between the real and the simulated point clouds. To enable the back-propagation from the feature loss to the previous raydrop module, we apply the gumble-softmax (Jang, Gu, and Poole 2016) on the ray drop processing.
@@
no raydrop random learning vgg rangenet
65.7
@@
Comparisons with State of the Art
CARLA and Real LiDAR In Table 1, we evaluate the quality of the generated data by using it to train 3D segmentation model (Zhou et al. 2021), the mean IoU is used as the evaluation metric.
@@
Reconstruction-based Simulators Our NeRF-LiDAR possesses apparent advantages over other reconstruction-based LiDAR simulators (Manivasagam et al. 2020; Fang et al. 2020). First, RGB images are used to assist the reconstruction in our NeRF-LiDAR, providing useful multi-view geometry information for more accurate reconstruction and label generation. Secondly, no manual annotations on the point clouds are required. We use NeRF representation to learn label generation. Multi-view images provide useful label information and geometry consistency to reduce the outlier labels. Moreover, no mesh reconstruction of the real world is required. NeRF-LiDAR can preserve more details of the real world by leveraging the RGB texture contents. Finally, NeRF-LiDAR does not require dense point clouds for the reconstruction of the real world. LiDARSim (Manivasagam et al. 2020) uses a 64-channel LiDAR and scans the street many times to achieve the dense point clouds for LiDAR simulation. Augmented LiDAR simulator (Fang et al. 2020) utilizes a special and expensive LiDAR device to achieve the dense point clouds. As a comparison, NeRF-LiDAR relies more on the 2D images to achieve 3D geometry information.
@@
Combining Real and Simulated Data We combine real data with NeRF-LiDAR data generated from ground-truth scenes to see if simulated data can further boost performance when used for training. As shown in Table 1, simulated data along with 10% real data (real 1k) are able to achieve better performance (Sim10k + real 1k: 89.1) than 100% real data (real 10k: 88.4).
@@ Conclusion
Acknowledgments
This work was supported in part by STI2030-Major Projects (Grant No. 2021ZD0200204), National Natural Science Foundation of China (Grant No. 62106050 and 62376060), Natural Science Foundation of Shanghai (Grant No. 22ZR1407500) and USyd-Fudan BISA Flagship Research Program.
@@ Appendix A More Experimental Results and Details
Table 5: Evaluation on the Test Set 1, where the simulation data and the test set is from the same driving scenes. Rangenet++(Milioto et al. 2019) is trained on the different training sets for evaluations. $*{}^{*}$ mean IoU of four classes.
@@ Appendix B More Comparisons and Analysis
Data Augmentation
Apart from the pure simulation comparison (Table 4 of the main paper), we have additionally provided ‘LiDARSim + Real 1k’ for consistent evaluation.
road vege. terrain vehicle manmade mIoU
Real 1k 96.2 83.6 83.1 83.0 86.4 86.5
LiDARSim + Real 1k 96.6 83.6 85.1 86.9 86.8 87.8
Ours + Real 1k 97.1 84.1 85.3 92.2 86.9 89.1
Table 13: Sim + Real experiments.